Solve the formula for the indicated varible. p = c + 2b + 2a, for c
Answer by algebrahouse.com(1659)

"Solve the formula for the indicated varible. p = c + 2b + 2a, for c"

p = c + 2b + 2a
c = p - 2b - 2a {subtracted 2b and 2a from both sides}

For more help from me, visit: www.algebrahouse.com